# Phase-Space Engineering and Dynamical Long-Range Order in Memcomputing

### v1

@@ IV Numerical Results
We are now ready to study in detail the relevance of the different physical hyper-parameters on the DMM’s ability to solve instances. Equations (11) through (16) feature five parameters: $\alpha$, $\beta$, $\gamma$, $\delta$, and $\zeta$. $\alpha$ and $\beta$ influence the rate of change of the short- and long-term memories, respectively (with $\alpha<\beta$). $\gamma$ and $\delta$ are thresholding parameters that must both be less than $1/2$. $\zeta$ influences the degree of “rigidity” in the system, or more precisely, the long-term memories’ relative influence on the system’s rigidity (see Ref. [9] for more discussion on the role of these quantities). In this study, we fix $\gamma$, $\delta$, and the relative value of $\alpha/\beta$. We then exclusively concentrate on the 2D parameter space $\{\beta,\zeta\}$, as these two parameters appear to influence the phase space dynamics most strongly.
@@ IV.1 Range of viable hyper-parameters
Figure 1: (a) State space diagram in hyper-parameter space $\{\beta,\zeta\}$, showing the maximum size 3-SAT ($\alpha_{r}=4.3$) instance $N_{max}$ that a DMM can solve within a fixed number of integration steps $T_{median}$ with fixed $dt_{0}=1.0$. The number of logical variables, $N$, is increased by $10$ at a time until $T_{median}$ exceeds $10^{6}$ steps. We find a wide region in phase space in which moderately large $N$ instances ($\sim 400-600$) can be solved quickly without tuning $dt_{0}$ (which scales the adaptive time step). Colored dots and stars correspond to parameter choices depicted in (b). (b) Scalability plots of the median number of steps to find a solution $T_{median}$ as a function of number of logical variables $N$. For two sets of parameters ($\{\beta,\zeta\}=\{\beta_{opt},\zeta_{opt}\}=\{79,0.063\}$ and $\{6.0\beta_{opt},1.5\zeta_{opt}\}$), the $T_{median}(N)$’s are fitted well by polynomials ($\sim N^{1.65}$ and $\sim N^{3.18}$, respectively). Outside the phase of “good” parameters, $T_{median}(N)$’s are fitted well by exponentials ($\sim e^{0.12N}$, $\sim e^{0.07N}$, and $\sim e^{0.04N}$ for $\{\beta,\zeta\}=\{\beta_{opt},\zeta_{opt}/10\}$, $\{\beta_{opt}/10,\zeta_{opt}\}$, and $\{10\beta_{opt},\zeta_{opt}\}$, respectively). Here, $dt_{0}$ is tuned at each data point to control numerical errors; everywhere, it takes a value between $0.01$ and $1.0$. In both figures, batches of $100$ instances are run in parallel, and the batch is considered solved once the median of its instances is solved. $\{\alpha/\beta,\beta,\gamma,\delta,\zeta\}_{opt}=\{0.45,79,0.36,0.080,0.063\}$ in both (a) and (b).
@@
IV.2 DLRO, but not necessarily scale-invariance
@@
Figure 2: (a) Diagram in $\{\beta,\zeta\}$ (in units of $\{\beta_{opt},\zeta_{opt}\}=\{79,0.063\}$) characterizing the degree of scale-freeness of avalanche distributions, alongside (b)-(e) four of these avalanche distributions at four points in parameter space, labeled in panel (a). In diagram (a), a parameter range exists where scale-free (SF) avalanche distributions are detected with an exponent between $2$ and $3$. Of the four points in the diagram (a) whose distributions are shown, three are not SF (points (b), (d), and (e)). However, some of these distributions (at points (b) and (e)) still feature large avalanches that approach the system size. At point (c), the distribution is fitted by $s^{-2.65}$ and includes a characteristic bump due to finite size effects. At point (d), only small avalanches are detected, as the lack of memory in the system destroys DLRO. 100 batches are simulated, each batch includes 100 instances each with $N=100$, and a run is stopped after $T=1000$ steps. The time window to detect the avalanches is $\Delta_{tw}=1.0$, $0.006$, $0.0006$, and $0.035$ for points (b), (c), (d), and (e), respectively (see the SM for more details).
@@
As anticipated in Sec. II, DLRO is the key physical property that enables DMMs to efficiently find a solution to a given problem. To corroborate this, we study the emergence of correlations in DMMs by extracting avalanche distributions of the fast DOFs. We define an avalanche as a cluster of connected “flips” in the logical variables that occur in recent succession (see the SM for a more detailed definition). These can be visualized as “waves” that propagate through the system comprised of many $v_{i}$’s whose corresponding Boolean values flip rapidly. In our phase space dynamics, avalanches correspond to jumps along instantons.
@@
In Fig. 2, we plot the associated scale-free exponents, if they exist, of these avalanche distributions at many points in our $\{\beta,\zeta\}$ parameter space. In a wide region roughly centered at $\{\beta_{opt},\zeta_{opt}\}=\{79,0.063\}$, we detect scale-free (SF) distributions with exponents between $2$ and $3$. One such distribution, at $\{\beta,\zeta\}=\{3\beta_{opt},\zeta_{opt}\}$ (point (c) in Fig. 2(a)), is plotted in panel Fig. 2(c).
@@
Interestingly, at most other points in parameter space, avalanche distributions still feature large (up to the system size $N$) avalanches. This is not too surprising since many parameters were already shown in Fig. 1(a) to be viable, and when $\beta$ is low, we expect parameters to be poor due to the prevalence of time-reversed trajectories (anti-instantons), not a lack of DLRO (cf. Sec. IV.4). Examples include $\{\beta,\zeta\}=\{0.08\beta_{opt},20\zeta_{opt}\}$ and $\{20\beta_{opt},20\zeta_{opt}\}$ (points (b) and (e) Fig. 2(a), respectively), both of which are plotted Fig. 2(b) and (e). This suggests that, even if DMMs do not possess perfect scale-invariance for some parameter choices, their memories are nonetheless quite effective at inducing long-range correlations in the logical variables. It is only due to the strong coupling between $v_{i}$’s (i.e., their collective behavior) due to memory, that large clusters of $v_{i}$’s can all “flip” their logical values in short succession, corresponding to such large avalanches.
@@
We suspect that, in the thermodynamic limit, the subset of viable parameters (the bright, central region in Fig. 1(a)) should converge to the parameter regime that yields SF avalanche distributions. As evidence, consider the distributions in Fig. 2(b) and (e). While they do feature large avalanches approaching size $N=10^{2}$, their correlations decay faster than the SF case at point (c) (Fig. 2(c)), and they showcase no finite size “bump”. Therefore, if we were to perform simulations with larger $N$ at points 1 and 4, we would not expect the size of the largest avalanches to increase much beyond $10^{2}$. As $N$ is increased further, arbitrarily large correlated clusters are required to solve instances efficiently, meaning SF-ness would become necessary in the large-$N$ limit. Unfortunately, our computational resources prevent us from extracting avalanche distributions for larger instances with reasonable statistics.
IV.3 Loss of DLRO when memories are too fast
@@
Crucially, point (d) ($\{\beta,\zeta\}=\{20\beta_{opt},0.08\zeta_{opt}\}$) in Fig. 2(d)) does not feature any large-scale avalanches. The reason is quite easy to understand. From our general analysis in Sec.II, we found that $\beta$ cannot be too large if DLRO is to be established, which is necessary for the DMM to capitalize on the global structure of the problem. Therefore, we anticipated that once $\beta$ exceeded a particular value, memory-induced DLRO should break down and the DMM would have to rely on purely “local” information.
@@
However, recall from that analysis that we needed $\beta\lesssim|\sum_{m}\overline{C_{m}^{(0)}(\vec{v})}\,\overline{A_{n,m}^{(1)}(\vec{v})}|$ for DLRO to emerge, which implicitly depends on $A_{n,m}(\vec{v})=x_{m}^{l}G_{n,m}(v_{n},v_{j},v_{k})-(1+\zeta x_{m}^{l})R_{n,m}(v_{n},v_{j},v_{k})$. Therefore, by varying $\zeta$, we also vary the timescale associated with the logical DOFs. This leads to the “stabilizing” effect present in Fig. 1(a); when $\beta$ and $\zeta$ are large simultaneously, timescale separation between the logical and memory DOFs, and thus DLRO, persists. This effect is less pronounced when $\beta$ and $\zeta$ are instead both small, as the term $(1+\zeta x_{m}^{l})R_{n,m}$ helps to bound the logical DOF timescale from above. This is precisely what we observe numerically; when $\beta$ is too large (and $\zeta$ does not accordingly increase to offset it), the timescale separation between the memory and logical DOFs is destroyed, eliminating DLRO. Here, it corresponds to a lack of large-scale avalanches.
@@ IV.4 Anti-instantons when memories are too slow
The corresponding time-reversed processes (also known as anti-instantons) correspond to a collection of flipping $v_{i}$’s which instead drive the system away from the solution (i.e., towards a saddle with larger index). Typically, this process is exponentially suppressed [3], but it can occur more often when the noise (whether physical or numerical) is large enough to overcome the “energetic” barrier for the system to transition from a critical point with given stability to a less stable one  [^7] This is similar to water flowing downhill: the water cannot flow uphill unless strong noise is introduced into the system..
@@
We plot the number of anti-instantons detected as a function of $\beta$ and $\zeta$ in Fig. 3. They only occur when $\beta$ is small. Small $\beta$ corresponds to very slowly varying memory DOFs, meaning there is a much lower degree of curvature in the “memory directions” in phase space. So, the numerical errors that accumulate in the solution search will be more problematic in this regime, since the separation of “energy levels” between saddles on either side of an instanton can become quite small  [^8] Note that an explicit calculation of the curvature of the stable and unstable manifolds near saddles points is computationally intractable for multi-dimensional phase spaces. Therefore, we find this method of extracting anti-instantons to be more efficient and sufficient evidence to describe the poor performance of DMMs in the small-$\beta$ regime.. Therefore, only a small amount of numerical noise is required to push the system against the flow that leads to the solution  [^9] In the analogy of water flowing downhill, it takes considerably less noise for the water to occasionally flow uphill, the smaller the slope of the hill. In the limit of zero slope, the water has the same probability of flowing “downhill” as “uphill” in the presence of noise..
@@ IV.5 Time-to-Solution Distributions
Interestingly, at all points in parameter space, even when parameters lead to poor performance, the TTS distributions are fitted reasonably well by IGs. The fact that distributions are still IGs even in this case, indicates that the DMMs’ characteristic phase space dynamics are largely unchanged. Both in the case of $\{\beta,\zeta\}=\{0.08\beta_{opt},20\zeta_{opt}\}$ (Fig. 4(a)), where we detect the most anti-instantons due to numerical noise, and $\{\beta,\zeta\}=\{20\beta_{opt},0.08\zeta_{opt}\}$ (Fig. 4(c)), where the lack of timescale separation destroys DLRO, the distributions have this same characteristic shape. This suggests that the DMM’s phase space topology remains unaffected even after varying parameters away from optimality by more than an order of magnitude.
In other words, although the lack of DLRO and the existence of anti-instantons certainly hinder a DMM’s solution search, neither seem to significantly push the system off the composite instanton manifold. When DLRO disappears, some instantonic trajectories (those corresponding to many logical DOFs flipping) become inaccessible, so a longer path in phase space must be taken to reach the equilibrium. When there are anti-instantons, the system may reverse its path along an instanton, nullifying some previous progress towards the equilibrium. Importantly, in neither case does the system need to navigate its entire $N$-dimensional phase space (ergodicity is still broken). This clarifies why these difficult instances can be solved at all even with such a poor choice of parameters (albeit with correspondingly poor scalability as a function of problem size).
@@ V Conclusions
DLRO in fast DOFs induced by memory (slow DOFs) has already been shown to exist in many other physical contexts, from spin systems [24] to neuromorphic devices [23] to models of human cortical dynamics [25]. DMMs are yet another example of dynamical systems with two distinct DOFs in which such a phenomenon emerges and is further leveraged to solve combinatorial optimization problems.
In this work, we have performed an extensive study of the role of the hyper-parameters entering the ODEs of DMMs to engineer the DMMs’ phase space and affect DLRO. We find a wide range of parameters to be “viable”, meaning that the DMMs with those parameters are able to reach equilibrium (solution of the problem to solve) quite efficiently.
When there is no longer sufficient timescale separation between the two DOFs, DLRO is suppressed and the solution search by DMMs becomes inefficient. A similar inefficiency occurs when the memory DOFs are extremely slow. However, in this case, the origin of such inefficiency is related to the emergence of (numerical) noise-induced time-reversed trajectories (anti-instantons), which occasionally reverse the flow toward the solution, hindering the search process. We expect a similar phenomenon to emerge for the hardware version of these machines when they are subject to physical noise.
Although only a subset of viable parameters provide scale invariance (as suggested by the avalanche distributions), the latter does not seem to be always necessary to solve problem instances up to the sizes we could reach with our computational resources. Many other regions in parameter space feature large avalanches up to the largest problem size, suggesting that strong coupling among the logical DOFs (and thus, DLRO) still exists even without such scale invariance. However, further (analytical) work is needed to understand if the phase of DLRO will eventually “merge” with the phase of scale invariance in the thermodynamic limit, and what the consequences are for the efficiency of DMMs to solve problem instances in such a limit.
In addition to shedding further light on the physics behind these machines, this work paves the way for more efficient methods to tune the relevant hyper-parameters of the ODEs of DMMs. In fact, we anticipate that such tuning procedures would have to rely on the maximization of the DLRO between the fast DOFs, which is the most important physical advantage of these machines. We leave this for future studies.

### v2

@@ IV Numerical results
We are now ready to study in detail the relevance of the different physical hyper-parameters on the DMM’s ability to solve instances. In this study, we fix $\gamma$, $\delta$, and the relative value of $\alpha/\beta$. We then exclusively concentrate on the 2D parameter space $\{\beta,\zeta\}$, as these two parameters appear to influence the phase space dynamics most strongly.
@@ IV.1 Range of viable hyper-parameters
To contextualize these results, we compare against three modern physics-inspired Ising solvers: coherent Ising machines [28], gain-dissipative dynamics [29], and simulated bifurcation machines [30]. These methods optimize an Ising (quadratic unconstrained binary optimization) objective and therefore require a SAT-to-Ising mapping [36, 37]. On the planted 3-SAT benchmarks considered here, they exhibit exponential time-to-solution scaling, and none were able to solve instances with $N>30$ within $10^{5}$ steps. Comparison with other SAT solvers, such as WalkSAT and Survey Inspired Decimation were carried out previously in Ref. [25], showing again DMMs’ superior success probabilities and scaling behavior across system sizes.
@@
Figure 2: (a) Diagram in $\{\beta,\zeta\}$ (in units of $\{\beta_{opt},\zeta_{opt}\}=\{79,0.063\}$) characterizing the degree of scale-freeness of avalanche distributions, alongside (b)-(e) four of these avalanche distributions at four points in parameter space, labeled in panel (a). In diagram (a), a parameter range exists where scale-free (SF) avalanche distributions are detected with an exponent between $2$ and $3$. Of the four points in the diagram (a) whose distributions are shown, three are not SF (points b, d, and e). However, some of these distributions (at points b and e) still feature large avalanches that approach the system size. At point c, the distribution is fitted by $s^{-2.82}$ and includes a characteristic bump due to finite size effects, with finite size scaling plotted in the inset. At point d, only small avalanches are detected, as the lack of memory in the system destroys the collective behavior. 100 batches are simulated, and each batch includes 100 instances run for $T=1000$ steps with $N=100$ (except for (c), where $N=60$ is plotted in orange, and $N=80$ is plotted in green alongside $N=100$ in blue). The time window to detect the avalanches is $\Delta_{tw}=1.0$, $0.006$, $0.0006$, and $0.035$ for points b, c, d, and e, respectively (see the SM for more details).
IV.2 Collective dynamics, but not necessarily scale-invariance
@@
As anticipated in Sec. II, collective behavior is the key physical property that enables DMMs to efficiently find a solution to a given problem. To corroborate this, we study the emergence of correlations in DMMs by extracting avalanche distributions of the fast DOFs. We define an avalanche as follows. When the boolean value of any $v_{i}$ changes, we say an event has occurred at time $t$ and location $i$. We then search for similar flipping events at nearest-neighbors of $v_{i}$, defined as the other logical variables that appear in the same clauses as $v_{i}$, within a window of time. If one of the nearest-neighbors of $v_{i}$ flips before this window of time has passed, the size of the avalanche is increased by 1, and further events are then searched for adjacent to the newest event. See the SM for more details regarding the generation of these avalanches. These can be visualized as “waves” that propagate through the system comprised of many $v_{i}$’s whose corresponding boolean values flip rapidly. In our phase space dynamics, avalanches correspond to jumps along instantons.
@@
In Fig. 2, we plot the associated scale-free exponents, if they exist, of these avalanche distributions at many points in our $\{\beta,\zeta\}$ parameter space. In a wide region roughly centered at $\{\beta_{opt},\zeta_{opt}\}=\{79,0.063\}$, we detect scale-free (SF) distributions with exponents between $2$ and $3$. One such distribution, at $\{\beta,\zeta\}=\{3\beta_{opt},\zeta_{opt}\}$ (point c in Fig. 2(a)), is plotted in panel Fig. 2(c). In its inset, we perform a finite size scaling analysis according to the finite size scaling ansatz $P(s)\sim s^{-\alpha^{\prime}}\exp(-s/N)$ [38], where $\alpha^{\prime}$ is the fitted power law exponent. The fact that the distributions at sizes $N=60$, $80$, and $100$ collapse after imposing this ansatz suggests these avalanche distributions’ deviations from power laws (in particular, the finite size “bumps”) can be attributed to finite size effects, and that these distributions do approach power laws in the thermodynamic limit.
@@
Interestingly, at most other points in parameter space, avalanche distributions still feature large (up to the system size $N$) avalanches. This is not too surprising since many parameters were already shown in Fig. 1(a) to be viable, and when $\beta$ is low, we expect parameters to be poor due to the prevalence of time-reversed trajectories (anti-instantons), not a lack of collective behavior (cf. Sec. IV.4). Examples include $\{\beta,\zeta\}=\{0.08\beta_{opt},20\zeta_{opt}\}$ and $\{20\beta_{opt},20\zeta_{opt}\}$ (points b and e in Fig. 2(a), respectively), both of which are plotted Fig. 2(b) and (e). This suggests that, even if DMMs do not possess perfect scale-invariance for some parameter choices, their memories are nonetheless quite effective at inducing system-wide correlations in the logical variables. It is only due to the strong coupling between $v_{i}$’s (i.e., their collective behavior) due to memory, that large clusters of $v_{i}$’s can all “flip” their logical values in short succession, corresponding to such large avalanches.
@@
We expect that, in the thermodynamic limit, the subset of viable parameters (the bright, central region in Fig. 1(a)) should converge to the parameter regime that yields SF avalanche distributions. As evidence, we note the direct overlap between the black and green dots in Fig. 1(a) (whose $T_{median}(N)$ distributions are fitted well by polynomials) and the region of scale-freeness in Fig. 2(a). Furthermore, consider the distributions in Fig. 2(b) and (e). While they do feature large avalanches approaching size $N=10^{2}$, their correlations decay faster than the SF case at point c (Fig. 2(c)), and they showcase no finite size “bump”. Therefore, if we were to perform simulations with larger $N$ at points b and e, we would not expect the size of the largest avalanches to increase much beyond $10^{2}$. As $N$ is increased further, arbitrarily large correlated clusters are required to solve instances efficiently, meaning SF-ness would become necessary in the large-$N$ limit. Unfortunately, our computational resources prevent us from extracting avalanche distributions for larger instances with reasonable statistics.
IV.3 Loss of collective behavior when memories are too fast
@@
Crucially, point d ($\{\beta,\zeta\}=\{20\beta_{opt},0.08\zeta_{opt}\}$) in Fig. 2(d)) does not feature any large-scale avalanches. The reason is quite easy to understand. From our general analysis in Sec. II, we found that $\beta$ cannot be too large if collective dynamics are to be established, which is necessary for the DMM to capitalize on the global structure of the problem. Therefore, we anticipated that once $\beta$ exceeded a particular value, memory-induced collective behavior should break down and the DMM would have to rely on purely “local” information.
@@
However, recall from that analysis that we needed $\beta\lesssim|\sum_{m}\overline{C_{m}^{(0)}(\vec{v})}\,\overline{A_{n,m}^{(1)}(\vec{v})}|$ in order for $v_{i}$ to evolve collectively, which implicitly depends on $A_{n,m}(\vec{v})=x_{m}^{l}G_{n,m}(v_{n},v_{j},v_{k})-(1+\zeta x_{m}^{l})R_{n,m}(v_{n},v_{j},v_{k})$. Therefore, by varying $\zeta$, we also vary the timescale associated with the logical DOFs. This leads to the “stabilizing” effect present in Fig. 1(a); when $\beta$ and $\zeta$ are large simultaneously, timescale separation between the logical and memory DOFs, and thus collective behavior, persists. This effect is less pronounced when $\beta$ and $\zeta$ are instead both small, as the term $(1+\zeta x_{m}^{l})R_{n,m}$ helps to bound the logical DOF timescale from above. This is precisely what we observe numerically; when $\beta$ is too large (and $\zeta$ does not accordingly increase to offset it), the timescale separation between the memory and logical DOFs is destroyed, eliminating the collective dynamics in the fast DOFs. Here, it corresponds to a lack of large-scale avalanches.
@@ IV.4 Anti-instantons when memories are too slow
The corresponding time-reversed processes (also known as anti-instantons) correspond to a collection of flipping $v_{i}$’s which instead drive the system away from the solution (i.e., towards a saddle with larger index). Typically, this process is exponentially suppressed [3], but it can occur more often when the noise (whether physical or numerical) is large enough to overcome the “energetic” barrier for the system to transition from a critical point with given stability to a less stable one. This is similar to water flowing downhill: the water cannot flow uphill unless strong noise is introduced into the system.
@@
We plot the number of anti-instantons detected as a function of $\beta$ and $\zeta$ in Fig. 3. They only occur when $\beta$ is small. Small $\beta$ corresponds to very slowly varying memory DOFs, meaning there is a much lower degree of curvature in the “memory directions” in phase space. So, the numerical errors that accumulate in the solution search will be more problematic in this regime, since the separation of “energy levels” between saddles on either side of an instanton can become quite small. Therefore, only a small amount of numerical noise is required to push the system against the flow that leads to the solution. In the analogy of water flowing downhill, it takes considerably less noise for the water to occasionally flow uphill, the smaller the slope of the hill. In the limit of zero slope, the water has the same probability of flowing “downhill” as “uphill” in the presence of noise.
Note that an explicit calculation of the curvature of the stable and unstable manifolds near saddles points is computationally intractable for multi-dimensional phase spaces. Therefore, we find this method of extracting anti-instantons to be more efficient and sufficient evidence to describe the poor performance of DMMs in the small-$\beta$ regime.
@@ IV.5 Time-to-solution distributions
Interestingly, at all points in parameter space, even when parameters lead to poor performance, the TTS distributions are fitted reasonably well by IGs. The fact that distributions are still IGs even in this case, indicates that the DMMs’ characteristic phase space dynamics are largely unchanged. Both in the case of $\{\beta,\zeta\}=\{0.08\beta_{opt},20\zeta_{opt}\}$ (Fig. 4(a)), where we detect the most anti-instantons due to numerical noise, and $\{\beta,\zeta\}=\{20\beta_{opt},0.08\zeta_{opt}\}$ (Fig. 4(c)), where the lack of timescale separation prevents the logical variables from evolving collectively, the distributions have this same characteristic shape. The distribution in Fig. 4(d) does deviate slightly from a pure inverse Gaussian, but we attribute this to the significant increase in phase space curvature at that point in parameter space, which is a purely geometric change to the phase space. This suggests that the DMM’s phase space topology remains unaffected even after varying parameters away from optimality by more than an order of magnitude.
In other words, although the lack of collective behavior and the existence of anti-instantons certainly hinder a DMM’s solution search, neither seem to significantly push the system off the composite instanton manifold. When there is no longer collective behavior in the fast DOFs, some instantonic trajectories (those corresponding to many logical DOFs flipping) become inaccessible, so a longer path in phase space must be taken to reach the equilibrium. When there are anti-instantons, the system may reverse its path along an instanton, nullifying some previous progress towards the equilibrium. Importantly, in neither case does the system need to navigate its entire $N$-dimensional phase space (ergodicity is still broken). This clarifies why these difficult instances can be solved at all even with such a poor choice of parameters (albeit with correspondingly poor scalability as a function of problem size).
@@ V Conclusions
Collective dynamics in fast DOFs induced by memory (slow DOFs) has already been shown to exist in many other physical contexts, from spin systems [7] to neuromorphic devices [10] to models of human cortical dynamics [11]. DMMs are yet another example of dynamical systems with two distinct DOFs in which such a phenomenon emerges and is further leveraged to solve combinatorial optimization problems.
In this work, we have performed an extensive study of the role of the hyper-parameters entering the ODEs of DMMs to engineer the DMMs’ phase space and affect the collective evolution of the logical variables. We find a wide range of parameters to be “viable”, meaning that the DMMs with those parameters are able to reach equilibrium (solution of the problem to solve) quite efficiently.
When there is no longer sufficient timescale separation between the two DOFs, the system no longer evolves collectively and the solution search by DMMs becomes inefficient. Practically, this corresponds to the hyper-parameter $\beta$ being too large (without $\zeta$ being increased accordingly, which would reintroduce timescale separation) so that the memory variables evolve too rapidly. A similar inefficiency occurs when the memory DOFs are extremely slow. However, in this case, the origin of such inefficiency is related to the emergence of (numerical) noise-induced time-reversed trajectories (anti-instantons), which occasionally reverse the flow toward the solution, hindering the search process. In our simulations, this occurs when $\beta$ is too small. We expect a similar phenomenon to emerge for the hardware version of these machines when they are subject to physical noise.
Although only a subset of viable parameters provide scale invariance (as suggested by the avalanche distributions), the latter does not seem to be always necessary to solve problem instances up to the sizes we could reach with our computational resources. Many other regions in parameter space feature large avalanches up to the largest problem size, suggesting that strong coupling among the logical DOFs (and thus, collective dynamics) still exists even without such scale invariance. However, further (analytical) work is needed to understand if the phase of collective behavior will eventually “merge” with the phase of scale invariance in the thermodynamic limit, and what the consequences are for the efficiency of DMMs to solve problem instances in such a limit.
In addition to shedding further light on the physics behind these machines, this work paves the way for more efficient methods to tune the relevant hyper-parameters of the ODEs of DMMs. In fact, we anticipate that such tuning procedures would have to rely on maximizing a metric which assesses the collective behavior in the fast DOFs, which is the most important physical advantage of these machines. We leave this for future studies.